Hello, can you help with the following formula
Given y=x^(1/(p-1))
Calculate P if y =2.65 and x = 0.745
Many thanks

Answer by stanbon(75887)

Take the log of both sides of the equation to get the "x" out of the exponent.
log y = (1/(p-1))log x
Then divide both sides by log x to get:
(logy)/(logx) = 1/(p-1)
Invert both sides to get:
p-1 = (logx)/(logy)
p = (logx)/(logy)-1
Substitute to get:
p= (log0.745)/(log2.65) -1
p=-1.30
Cheers,
Stan H.